Solar Designs is considering an investment in an expanded product line. Two possible types of expansion are being considered. After investigating the possible outcomes the company made the estimates shown in the following table:


Expansion A Expansion B
Initial investment $12,000 $12,000
Annual rate of return
• Pessimistic 16% 10%
• Most likely 20% 20%
• Optimistic 24% 30%


1. Determine the range of the rates of return for each of the two projects.
2. Which project is less risky? Why?
3. If you were making the investment decision, which one would you choose? Why? What does this imply about your feelings toward risk?
4. Assume that expansion B’s most likely outcome is 21% per year and that all other facts remain the same. Does this change your answer to part c? Why?
Answer #1

Expension A = 12000-7200/12000*100%=40%

Expension B=12000-7200/12000*100%=40%

source: Financial management textbook
answered by: calynchua96
Answer #2

a)range for A =24-16 =8% , range for B = 30-10 =20%


b)second project is more risky as it has wider range. i.e as it has more standard deviation


c)I will choose first investment as it is having lesser risk.I am risk averse person


d)no it will not change my decission in part c . as it is not affecting range
Answer #3
a. Assuming optimistic A is 40%, not 4%, the ranges are 16-40% and 10-30% b. Project A is less risky, better downside and upside. c. I would choose A. Higher returns on the upside, better returns on the downside. Same returns most likely. Same investment level. d. No, it does not. Likely is similar, and and everything else is better.